# Noise2Ghost: Self-supervised deep convolutional reconstruction for ghost imaging

### v1

@@ Abstract
Index Terms:
Ghost imaging, deep learning, image reconstruction, inverse problems, reconstruction algorithms.
I Introduction
Ghost imaging (GI) is a paradigm-changing technique that is mostly known and anticipated for its promise to cut radiation doses. However, these possible radiation dose advantages rely on a much more fundamental concept: The ability to choose the trade-off between spatial resolution, field-of-view (FoV), and acquisition time somewhat independently of the incident beam size. Traditionally, to spatially resolve diffused signals like x-ray fluorescence (XRF), one had to raster scan the object of interest with focused illumination (pencil-beam, PB, Fig. 1a). The PB focal spot size determines the spatial resolution. GI achieves that by probing extended regions of a sample at once with structured illumination (Fig. 1b). This provides spatially resolved information, whose resolution is independent of the beam size and detector pixel size, but depends instead on the size of the beam structures.
@@ I Introduction
Thanks to all these advantages, GI has the potential to spark many breakthroughs in applications in various fields ranging from biomedical imaging to remote sensing. However, GI has only seen limited applicability so far, especially in the above-mentioned nano-scale imaging applications. In these cases, the observed signals are affected by overwhelmingly high Poisson noise levels due to their very small photon fluxes. This noise type does not enjoy the intrinsic noise reduction of GI, known as Felgett’s advantage, thus reducing GI’s advantage over PB [7]. Here, we propose a GI reconstruction method particularly suited for working with acquisitions affected by high noise levels, regardless of whether the noise type is Gaussian or Poisson. Our method leverages the latest developments in unsupervised machine learning (ML) techniques, so it does not require high-quality reference data. This work is a step towards practical applications of GI, particularly for photon-limited scenarios like nano-scale emission imaging.
II Method
In this section, we present the proposed method Noise2Ghost (N2G). First, we introduce related methods. Then, we describe the signal formation (forward) model and present our proposed reconstruction method. Finally, we show how this method addresses random acquisition noise.
II-A Existing methods
@@
Self-supervised methods like Noise2Noise (N2N) use the same training procedure as the supervised methods, but they use noisy images as targets [17]. They only require two noisy realizations of the same measurement (e.g., two noisy images of the same scene) and exploit the fact that the only difference between the two measurements should be the noise. Thus, given one of the two measurements, they train the CNN to predict the other and vice versa. As previously mentioned, CNNs can capture small and large-scale correlations, leading them to learn the features of the measurements while discarding their uncorrelated parts, i.e., the noise. N2N can be adapted to work with inverse problems like computed tomography [18]. However, as we will also see in section II-E, it does not handle the so-called missing realization problem, which refers to artifacts arising from not having enough realizations to represent the imaged objects, and that is usually an important source of noise in GI reconstructions.
II-B Forward model and variational methods
@@
Unsupervised generative machine learning methods have taken inspiration from Eq. 2 [12], and developed an image recovery scheme that seeks to solve a very similar minimization problem, but on the output of a machine-learning model (typically a CNN), as follows:
@@
where $r\in\mathbb{R}^{N}$ is any input image, which could be randomly chosen, $N_{\theta}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ the model and $\theta\in\mathbb{R}^{T}$ is its parameters vector with $T$ different parameters. In [14], $r=W^{\dagger}y$ where $W^{\dagger}$ is the pseudo-inverse of $W$, making $r$ the least-squares reconstruction. The model $N_{\theta}$ learns to produce the image $\hat{x}$ that satisfies both the forward model Eq. 1 and the regularization $R$, weighted by the $\lambda$ parameter. Neural networks have notions of non-locality and non-linearity in their response, which can behave as an additional prior on top of $R$ [9] when the training process is stopped before convergence.
While Eq. 3 has proven to help with the reconstruction of under-determined acquisitions [14], it is not tuned to deal with strong random noise. Following the same derivation steps as in [18], the expected prediction error of the data fitting term in Eq. 3a is:
@@
II-C Proposed method: Noise2Ghost (N2G)
@@
Figure 2: Schematic representation of the proposed method: (2a) The partitioning of the realizations set, generating the partial reconstructions (sub-recs). (2b) The training procedure, which adjusts the model weights to minimize the residual between the projection of each partial prediction against the set of realizations not used for the its corresponding input. (2c) The prediction of the final reconstruction obtained from the averaged sub-reconstructions.
Here, we propose our CNN-based noise-tolerant unsupervised GI reconstruction method. It is inspired by DIP and N2N while addressing their limitations, namely accounting for the random noise in the optimization technique and the missing realization artifacts, respectively. We start from the assumption that by selecting a subset of the original realization set (i.e., the collection of all the acquired signals paired with the corresponding illumination patterns), we should obtain a degraded reconstruction due to a decrease in captured information. However, the degraded reconstruction should still represent the same signal as the reconstruction from the full realization set. This means that if we partition the realizations into subsets with equal numbers of realizations, we should obtain equivalent but different and degraded sub-reconstructions.
We partition the realizations in $k$ splits $k\in[1,K]$, producing $K$ different sub-reconstructions (i.e., a k-tuple of reconstructions). Each sub-reconstruction in the said k-tuple has unique noise characteristics compared to the other reconstructions, and it provides different information on the FoV. From the partitioned data vectors $y_{k}$ we compute the sub-reconstructions $x_{k}$ as follows (Fig. 2a):
@@
where $W_{k}^{\dagger}$ is pseudo-inverse of the partition forward model $W_{k}$, and $b_{k}$ is the clean data.
Each sub-reconstruction $x_{k}$ represents the same signal but has a different noise. In N2G, we feed each of these sub-reconstructions to the model, and then we optimize the following minimization problem (Fig. 2b) by comparing them against the realizations that were not used to reconstruct them:
@@
Since the different sub-reconstructions $x_{k}$ should represent the same signal, we would like the model to produce the same result for each. Thus, the model should learn to differentiate between the actual signal and random noise. This is in opposition to Eq. 3a, where the same realizations were used for the input reconstruction and as targets in the learning loss. Finally, Fig. 2c represents how Eq. 6b produces the reconstructed images.
II-D Self-supervised random noise reduction
We now derive the noise-reduction properties of Eq. 6a. Reconstruction noise comes from two main sources: Measurement noise; and the lack of information from a low number of realizations, which creates the so-called missing realization artifacts. These noise sources can be represented in the sub-reconstructions $x_{k}$ as follows:
@@ II-D Self-supervised random noise reduction
where $W_{k}^{\dagger}\epsilon_{k}=t_{k}$ is the pseudo-inverse of the bucket measurement noise, while the vector $v_{k}\subseteq\ker W_{k}$ describes spurious solutions that could be added to the reconstruction, and still fit the data. Our method focuses on the random measurement noise, and the desired noise reduction effect is given by the specific construct of the data fitting term from Eq. 6a. When projecting the predicted image onto the target data from the other splits, the projected noise from the $k1$ split will be inconsistent with the random noise from the other $K-1$ splits. This can be seen by substituting Eq. 7a and Eq. 5 into the data fitting term of Eq. 6a:
@@
As $\epsilon_{k}$ is independent of any other $\epsilon_{i}$ and their domains do not overlap (Appendix A from [19]), the expected prediction error of Eq. 8 becomes:
@@
where the term $\lVert W_{i}N_{\theta}(x_{k})-b_{i}\rVert_{2}^{2}$ is the supervised reconstruction error, and the term $\lVert\epsilon_{i}\rVert_{2}^{2}$ is the signal noise variance. In the supervised reconstruction term, the acquisition noise is not present. This means that it is equivalent to having noiseless realizations as the targets of our N2G reconstruction, thus providing the noise suppression characteristics of Noise2Ghost. The signal noise variance is just a constant in the objective function, which only depends on the acquisition noise and should not contribute to the objective function gradient. In reality, it may introduce some noise in the optimization gradients at very high noise levels. This could interfere with the reconstruction, guiding it towards local minima.
II-E Relationship with Noise2Inverse
The Noise2Inverse method addresses the problem in Eq. 6 directly in the reconstruction domain [19], by solving the following problem:
@@ II-E Relationship with Noise2Inverse
where $x_{i}=W_{i}^{\dagger}y_{i}$ is the pseudo-inverse (LS reconstruction) of the partial set of masks $W_{i}$ and corresponding buckets $y_{i}$ with $i\in[1,K]$. Theoretical grounds support this formulation, and it should retrieve the same result as Eq. 6, regarding its noise reduction aspect. However, it does not address the missing realization artifacts and suffers from strong technical limitations. In ghost imaging, artifacts arising from both missing realization noise and random noise are structured and have long-range correlations, i.e., they generally cover the whole FoV. Commonly used models like CNNs have a limited receptive field (i.e., the region around each pixel that the CNN probes and that provides the context for the said pixel) that may not extend to the entire FoV. Depending on the model, this could have a strong negative impact on the denoising performance.
More precisely, about the missing realization artifacts, in [19] it is stated that Eq. 11 would be unable to cope with them. The LS sub-reconstructions cope with the missing realizations by acting as if their bucket values were set to 0. This fixes one given solution $v_{k}$ for each $x_{k}$, out of the infinite solutions that satisfy Eq. 1. This specific choice of $v_{k}$ is likely incorrect, but it is used as the model learning target in Eq. 11. The formulation we propose in (Eq. 6) circumvents this limitation by using the measured buckets as “target”, and by leaving the freedom to the model to find a solution that satisfies the forward model in Eq. 1 in the reconstruction space.
A possible approach to overcome the just-discussed limitations of Eq. 11 would be to use variational methods (like in Eq. 2) to compute the sub-reconstructions. These non-linear methods are known to remove or strongly reduce these long-range artifacts. However, as discussed in [19], their non-linear nature contrasts with the assumptions of Eq. 11. This may then result in a degraded reconstruction quality.
II-F Data augmentation
Thanks to the assumption that any permutation (reordering) of the acquired realizations provides the same reconstruction, we can achieve simple and effective data augmentation: We can increase the number of training examples for the self-supervised method from the measured data. We permute our dataset in $p$ arrangements with $p\in[1,P]$ and partition each arrangement in $k$ splits $k\in[1,K]$, producing $KP$ different sub-reconstructions (i.e., $P$ k-tuples of reconstructions). In practice, the permutations $P$ provide unique mixtures of missing realization artifacts and random noise, while the $K$ splits are still the underlying mechanism for noise reduction. This modifies Eq. 6 into:
@@ II-F Data augmentation
II-G Models
We briefly discuss the machine learning models, in particular neural networks (NN), that N2G could leverage. The model should accept multiple images of the same object as input. This is incompatible with INRs, which learn coordinate-based representations of the signal and thus take coordinates as inputs. CNNs and other MLP-based NNs are instead good candidates. Both NN types take images as input and can learn short- and long-range pixel correlations. Thus, N2G is not tied to a specific NN architecture and can take advantage of future developments in that area. MLP-based NNs examples include the Visual Transformer (ViT) [20] and the swin-transformer [21]. CNNs examples include U-Net [11], DnCNN [22], and MS-D net [23]. Regarding the model size, one should try to work with small NNs to minimize the computation time and resources and to limit overfitting. For instance, we suggest models with parameter numbers in the order of 100k for a $100\times 100$ pixels image. However, some models like MS-D net could achieve similar results with 10-20k parameters for similar image sizes. In our tests, we had the best results with U-Net and DnCNN with around 100-300k parameters. Further details will be provided with the results.
III Results
@@
III-A Synthetic data generation & quantification
@@
We extracted a $100\times 100$ pixels image to be our ground truth (phantom, e.g., from Figs. 5f and 6f) from [24], a curated collection of chromosome images. We generated synthetic random masks with a normalized half-Gaussian distribution and forward projected the phantom to create the corresponding buckets. As discussed in section -B, we added Poisson noise to the buckets by first multiplying them by the maximum expected emitted number of photons per pixel. We fed the resulting mean expected photons per realizations to the Poisson distribution function. We then normalized the signals by dividing the noisy buckets by the same number of expected photons, resulting in signals:
@@
We use the constant $C$ to modulate the noise and evaluate its effect on the reconstruction performance of the proposed method.
We evaluate the reconstruction algorithms both visually and with multiple objective metrics: the mean square error (MSE), the peak signal-to-noise ratio (PSNR), the structural similarity index (SSIM), and the resolution, which is estimated from the Fourier ring correlation (FRC). Their formulas are provided in section -A, while here we shortly describe their meaning. The MSE evaluates the average pixel-wise difference between the phantom and the reconstruction intensity values. The SSIM calculates the similarity between these two images by comparing the inter-dependencies of spatially neighboring pixels over various window sizes [25]. PSNR is based on MSE, but it focuses on the peak intensity power instead of its average. Finally, the FRC computes the similarity in Fourier space between the phantom and the reconstruction for all the sampled frequencies. We estimate the resolution of the reconstruction by finding the intersection of the resulting curve with the threshold function defined in [26].
III-B Synthetic data reconstructions across various noise levels

### v2

@@ Abstract
journal: opticajournal
1 Introduction
Ghost imaging (GI) is a paradigm-changing technique that is mostly known and anticipated for its promise to cut radiation doses. However, these possible radiation dose advantages rely on a much more fundamental concept: The ability to choose the trade-off between spatial resolution, field-of-view (FoV), and acquisition time somewhat independently of the incident beam size. Traditionally, to spatially resolve diffused signals like x-ray fluorescence (XRF), one had to raster scan the object of interest with focused illumination (pencil-beam, PB, Fig.˜1a). The PB focal spot size determines the spatial resolution. Classical GI achieves that by probing extended regions of a sample at once with structured illumination (Fig.˜1b). This provides spatially resolved information, whose resolution is independent of the beam size and detector pixel size, but depends instead on the size of the beam structures. Hereafter, we will refer to classical GI simply as GI, which, unlike quantum GI [1], does not require pairs of quantum entangled photons.
@@ 1 Introduction
Thanks to all these advantages, GI has the potential to spark many breakthroughs in applications in various fields ranging from biomedical imaging to remote sensing. However, GI has only seen limited applicability so far, especially in the above-mentioned nano-scale imaging applications. In these cases, the observed signals are affected by overwhelmingly high Poisson noise levels due to their very small photon fluxes. This noise type does not enjoy the intrinsic noise reduction of GI, known as Felgett’s advantage, thus reducing GI’s advantage over PB [8]. Here, we propose a GI reconstruction method particularly suited for working with acquisitions affected by high noise levels, regardless of whether the noise type is Gaussian or Poisson. Our method leverages the latest developments in unsupervised machine learning (ML) techniques, so it does not require high-quality reference data. This work is a step towards practical applications of GI, particularly for photon-limited scenarios like nano-scale x-ray emission imaging [10].
2 Method
In this section, we present the proposed method, Noise2Ghost (N2G). First, we introduce related methods. Then, we describe the signal formation (forward) model and present our proposed reconstruction method. Finally, we show how this method addresses random acquisition noise. The code of the algorithms presented in this section is available at [11].
2.1 Existing methods
@@
Self-supervised methods like Noise2Noise (N2N) use the same training procedure as the supervised methods, but they use noisy images as targets [19]. They only require two noisy realizations of the same measurement (e.g., two noisy images of the same scene) and exploit the fact that the only difference between the two measurements should be the noise. Thus, given one of the two measurements, they train the CNN to predict the other and vice versa. As previously mentioned, CNNs can capture small and large-scale correlations, leading them to learn the features of the measurements while discarding their uncorrelated parts, i.e., the noise. N2N can be adapted to work with inverse problems like computed tomography [20]. However, as we will also see in Subsec.˜2.5, it does not handle the so-called missing realization problem, which refers to artifacts arising from not having enough realizations to represent the imaged objects, and that is usually an important source of noise in GI reconstructions.
2.2 Forward model and variational methods
@@
Unsupervised generative machine learning methods have taken inspiration from Eq.˜2 [15], and developed an image recovery scheme that seeks to solve a very similar minimization problem, but on the output of a machine-learning model (typically a CNN), as follows:
@@
where $r\in\mathbb{R}^{N}$ is any input image, which could be randomly chosen, $N_{\theta}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ the model, and $\theta\in\mathbb{R}^{T}$ is its parameters vector with $T$ different parameters. In [17], $r=W^{\dagger}y$ where $W^{\dagger}$ is the pseudo-inverse of $W$, making $r$ the least-squares reconstruction. The model $N_{\theta}$ learns to produce the image $\hat{x}$ that satisfies both the forward model Eq.˜1 and the regularization $R$, weighted by the $\lambda$ parameter. Neural networks have notions of non-locality and non-linearity in their response, which can behave as an additional prior on top of $R$ [12] when the training process is stopped before convergence. The $\lambda$ parameter is commonly determined through cross-validation, by creating a small leave-out set of measurements, and choosing the $\lambda$ that makes the model fit them best [21].
While Eq.˜3 has proven to help with the reconstruction of under-determined acquisitions [17], it is not tuned to deal with strong random noise. Following the same derivation steps as in [20], the expected prediction error of the data fitting term in Eq.˜3a is:
@@
2.3 Proposed method: Noise2Ghost (N2G)
@@
Figure 2: Schematic representation of the proposed method: (2a) The partitioning of the realizations set, generating the partial reconstructions (sub-recs), (2b) the training procedure, and (2c) the prediction of the final reconstruction.
Here, we propose our CNN-based noise-tolerant unsupervised GI reconstruction method. It is inspired by DIP and N2N while addressing their limitations, namely accounting for the random noise in the optimization technique and the missing realization artifacts, respectively. The former is addressed in Subsec.˜2.4, and the latter in Subsec.˜2.5. We start from the assumption that by selecting a subset of the original realization set (i.e., the collection of all the acquired signals paired with the corresponding illumination patterns), we should obtain a degraded reconstruction due to a decrease in captured information. However, the degraded reconstruction should still represent the same signal as the reconstruction from the full realization set. This means that if we partition the realizations into subsets with equal numbers of realizations, we should obtain equivalent but different and degraded sub-reconstructions.
We partition the realizations in $k$ splits $k\in[1,K]$, producing $K$ different sub-reconstructions (i.e., a k-tuple of reconstructions). Each sub-reconstruction in the said k-tuple has unique noise characteristics compared to the other reconstructions, and it provides different information on the FoV. From the partitioned data vectors $y_{k}$ we compute the sub-reconstructions $x_{k}$ as follows (Fig.˜2a):
@@
where $W_{k}^{\dagger}$ is pseudo-inverse of the partition forward model $W_{k}$, and $b_{k}$ is the clean data. The $x_{k}$ solutions can be reconstructed with a common least-squares solver.
Each sub-reconstruction $x_{k}$ represents the same signal but has a different noise. In N2G, we feed each of these sub-reconstructions to the model, and then we optimize the following minimization problem (Fig.˜2b) by comparing them against the realizations that were not used to reconstruct them:
@@
Since the different sub-reconstructions $x_{k}$ should represent the same signal, we would like the model to produce the same result for each. Thus, the model should learn to differentiate between the actual signal and random noise. This is in opposition to Eq.˜3a, where the same realizations were used for the input reconstruction and as targets in the learning loss. Finally, Fig.˜2c represents how Eq.˜6b produces the reconstructed images. In Sec. S4 of the Supplementary Information We briefly discuss the machine learning models, in particular neural networks (NN), that N2G could leverage.
2.4 Self-supervised random noise reduction
We now derive the noise-reduction properties of Eq.˜6a. Reconstruction noise comes from two main sources: Measurement noise; and the lack of information from a low number of realizations, which creates the so-called missing realization artifacts. These noise sources can be represented in the sub-reconstructions $x_{k}$ as follows:
@@ 2.4 Self-supervised random noise reduction
where $W_{k}^{\dagger}\epsilon_{k}=t_{k}$ is the pseudo-inverse of the bucket measurement noise, while the vector $v_{k}\subseteq\ker W_{k}$ describes spurious solutions that could be added to the reconstruction, and still fit the data. Our method focuses on the random measurement noise, and the desired noise reduction effect is given by the specific construct of the data fitting term from Eq.˜6a. When projecting the predicted image onto the target data from the other splits, the projected noise from the $k1$ split will be inconsistent with the random noise from the other K1K-1 splits. This can be seen by substituting Eq.˜7a and Eq.˜5 into the data fitting term of Eq.˜6a:
@@
As $\epsilon_{k}$ is independent of any other $\epsilon_{i}$ and their domains do not overlap (Appendix A from [22]), the expected prediction error of Eq.˜8 becomes:
@@
where the term $\lVert W_{i}N_{\theta}(x_{k})-b_{i}\rVert_{2}^{2}$ is the supervised reconstruction error, and the term $\lVert\epsilon_{i}\rVert_{2}^{2}$ is the signal noise variance. In the supervised reconstruction term, the acquisition noise is not present. This means that it is equivalent to having noiseless realizations as the targets of our N2G reconstruction, thus providing the noise suppression characteristics of Noise2Ghost. The signal noise variance is just a constant in the objective function, which only depends on the acquisition noise and should not contribute to the objective function gradient. In reality, it may introduce some noise in the optimization gradients at very high noise levels. This could interfere with the reconstruction, guiding it towards local minima. As for N2N and derived methods, this often results in excessive blur and resolution loss in the reconstructions.
2.5 Relationship with Noise2Inverse
The Noise2Inverse method addresses the problem in Eq.˜6 directly in the reconstruction domain [22], by solving the following problem:
@@ 2.5 Relationship with Noise2Inverse
where $x_{i}=W_{i}^{\dagger}y_{i}$ is the pseudo-inverse (LS reconstruction) of the partial set of masks $W_{i}$ and corresponding buckets $y_{i}$ with $i\in[1,K]$. Theoretical grounds support this formulation, and it should retrieve the same result as Eq.˜6, regarding its noise reduction aspect. However, it does not address the missing realization artifacts and suffers from strong technical limitations. In ghost imaging, artifacts arising from both missing realization noise and random noise are structured and have long-range correlations, i.e., they generally cover the whole FoV. Commonly used models like CNNs have a limited receptive field (i.e., the region around each pixel that the CNN probes and that provides the context for the said pixel) that may not extend to the entire FoV. Depending on the model, this could have a strong negative impact on the denoising performance.
More precisely, about the missing realization artifacts, in [22] it is stated that Eq.˜11 is unable to cope with them. The LS sub-reconstructions cope with the missing realizations by acting as if their bucket values were set to 0. This selects one specific solution $v_{k}$ for each $x_{k}$, out of the infinite solutions in $\ker W_{k}$ that satisfy Eq.˜1. This specific choice of $v_{k}$ is probably incorrect, but is imposed by N2I as the learning target in Eq.˜11. The model learns to discard the inconsistent features of different $v_{k}$ functions, selecting a specific $v\in\ker W$ as a common ground.
The formulation we propose in (Eq.˜6) circumvents this limitation by using the measured buckets as “target”. Thus, N2G does not embed any specific $v_{k}$ into the learning target. This allows the model to select a $v\in\ker W$ that satisfies both the measured buckets and the prior knowledge injected by the regularization term $R$ and the regularization properties of CNNs [12].
A possible approach to overcome the just-discussed limitations of Eq.˜11 would be to use variational methods (like in Eq.˜2) to compute the sub-reconstructions. These non-linear methods are known to remove or strongly reduce these long-range artifacts. However, as discussed in [22], their non-linear nature contrasts with the assumptions of Eq.˜11. This may then result in a degraded reconstruction quality.
2.6 Data augmentation
Thanks to the assumption that any permutation (reordering) of the acquired realizations provides the same reconstruction, we can achieve simple and effective data augmentation: We can increase the number of training examples for the self-supervised method from the measured data. We permute our dataset in $p$ arrangements with $p\in[1,P]$ and partition each arrangement in $k$ splits $k\in[1,K]$, producing $KP$ different sub-reconstructions (i.e., $P$ k-tuples of reconstructions). In practice, the permutations $P$ provide unique mixtures of missing realization artifacts and random noise, while the $K$ splits are still the underlying mechanism for noise reduction. This modifies Eq.˜6 into:
@@ 2.6 Data augmentation
3 Results
@@
3.1 Synthetic data generation & quantification
@@
We extracted a $100\times 100$ pixels image to be our ground truth (phantom, e.g., from Figs.˜5f and 6f) from [23], a curated collection of chromosome images. We generated synthetic random masks with a normalized half-Gaussian distribution and forward projected the phantom to create the corresponding buckets. We added Poisson noise to the buckets by first multiplying them by the maximum expected emitted number of photons per pixel. We fed the resulting mean expected photons per realizations to the Poisson distribution function. We then normalized the signals by dividing the noisy buckets by the same number of expected photons, resulting in signals:
@@
This expression is derived in Sec. 2 of the Supplementary Information for the specific case of XRF imaging. We use the constant $C$ to modulate the noise and evaluate its effect on the reconstruction performance of the proposed method. The advantage of this noise model is that it gives emphasis to the number of photons delivered to the sample.
We evaluate the reconstruction algorithms both visually and with multiple objective metrics: the mean square error (MSE), the peak signal-to-noise ratio (PSNR), the structural similarity index (SSIM), and the resolution, which is estimated from the Fourier ring correlation (FRC). Their formulas are provided in Sec. 1 of the Supplementary Information, while here we shortly describe their meaning. The MSE evaluates the average pixel-wise difference between the phantom and the reconstruction intensity values. The SSIM calculates the similarity between these two images by comparing the inter-dependencies of spatially neighboring pixels over various window sizes [24]. PSNR is based on MSE, but it focuses on the peak intensity power instead of its average. Finally, the FRC computes the similarity in Fourier space between the phantom and the reconstruction for all the sampled frequencies. We estimate the resolution of the reconstruction by finding the intersection of the resulting curve with the threshold function defined in [25].
3.2 Synthetic data reconstructions across various noise levels